OpenStudy (anonymous):

3√-27 step by step please thanks

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

first find the root of -27

OpenStudy (anonymous):

it should be \[3\sqrt{-3}\]

OpenStudy (anonymous):

then multiply the three

OpenStudy (anonymous):

-3 ?? O_O

OpenStudy (anonymous):

and get \[9\sqrt{-3}\] iguess

OpenStudy (anonymous):

yes radicals are confusing

OpenStudy (anonymous):

\[3\sqrt{-27}\] then, the largest square into that is 9, aka \[3\sqrt{3\times3\times-3}\] then take the two 3s out, and you get\[3\times3\sqrt{-3}\] add together like terms and you get\[9\sqrt{-3}\] BUT you cant ever have a negative radical, so you multiply by i, which is an imiganary number for radicals (basically multiplying by -1) and in the end you get \[9i \sqrt{3}\] which is your final answer

OpenStudy (anonymous):

oh, well guessed i learned something too

OpenStudy (anonymous):

im a little further ahead in math i guess, so this is kinda simple, but a good base point you should know

OpenStudy (anonymous):

I just hate square roots -.-